Find: (6/13) ÷ 7 Find: (6/13) ÷ 7 Find: (6/13) ÷ 7 Find: (6/13) ÷ 7

 
Find:
  (6/13) ÷ 7
Find:
  (6/13) ÷ 7
Find:
  (6/13) ÷ 7
Find:
  (6/13) ÷ 7

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= (6/13) × reciprocal of 7
= (6/13) × (1/7)
= (6 × 1) / (13 × 7)
= 6/91

Thanks